OpenStudy (anonymous):

1. A department store pays $20 for each watch in a shipment. The store sells the watches for $75. What is the maximum percent of discount the store can offer and make a profit of at least $20 on each watch? Round your answer to the nearest hundredth. (1 point)

OpenStudy (anonymous):

If the profit is AT LEAST $20, then the SP should be $40. So maximum discount = 35/75*100 = 7/15 * 100 = 7/3 * 20 = 140/3 = 46 2/3%

OpenStudy (anonymous):

is 46.67 the answer

OpenStudy (anonymous):

which is 46 2/3

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Maximum selling price is equal to $20 + $20 = $40. 1% off $75 is $0.75 $75-$40 = $35 maximum discount. 35/0.75 = 46.67 => 47%

OpenStudy (anonymous):

approximately 47%.